4 Find the general solution for y – 3y +2y = 1+e-?

Detailed answer, showing the integration
Answer #1

y- 3yltzy. Het (1?-35+2)y= 14 Het of the form f1D).J= O(t) where fib) = 1²3p+2. Out) Auxilary equation as {(m)=0 m?_gmt2=0 -.t -4 Het Put q=et dq=-etde put Patet dp =-etat. 497 dq +q -- ef de - 4 Jeg et th) in na da 4 loge q+)--) da 1+q 4 log (itet)